An object of mass 4kg moves round a circle of radius 6m with a constant speed of 12m/s. Calculate (i) the angular speed (ii) the force towards the center.
1
Expert's answer
2020-11-25T10:54:36-0500

Answer

Angular speed

"\omega =\frac{v}{r}\\=\frac{12}{6}\\=2rad/s"


Force towards center

"F=\frac{mv^2}{r}\\F=\frac{4\times12^2}{6}\\F=96N"